OpenStudy (anonymous):

What is the equation of the line that passes through (0, -2) and has a slope of 0? a) y=-2 b) y=2 c) x=2

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\] m=slope (x1,y1)

OpenStudy (pawanyadav):

Only you have to do is putting values in y=mx+c

pooja195 (pooja195):

@Pawanyadav you can give the user some time rather than being rude.

OpenStudy (anonymous):

I had missed read it.. and I read it again and got it